Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows: Standard Quantity or Hours Standard Price or Rate Standard Cost Direct materials 2.30 ounces $ 26.00 per ounce $ 59.80 Direct labor 0.50 hours $ 14.00 per hour 7.00 Variable manufacturing overhead 0.50 hours $ 3.40 per hour 1.70 Total standard cost per unit $ 68.50 During November, the following activity was recorded related to the production of Fludex: Materials purchased, 12,500 ounces at a cost of $305,625. There was no beginning inventory of materials; however, at the end of the month, 2,800 ounces of material remained in ending inventory. The company employs 21 lab technicians to work on the production of Fludex. During November, they each worked an average of 150 hours at an average pay rate of $12.00 per hour. Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs during November totaled $4,200. During November, the company produced 4,200 units of Fludex. Required: 1. For direct materials: a. Compute the price and quantity variances. b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract? 2. For direct labor: a. Compute the rate and efficiency variances. b. In the past, the 21 technicians employed in the production of Fludex consisted of 4 senior technicians and 17 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued? 3. Compute the variable overhead rate and efficiency variances.
Answer #1

1.

Direct material price variance = Actual quantity*standard price - Actual quantity*actual price

Direct material price variance = 12,500*$26 - $305,625

Direct material price variance = $325,000 - 305,625 = $19,375 Favorable

Direct material quantity variance = Standard quantity*standard price - Actual quantity*standard price

Direct material quantity variance = (4,200*2.30)*$26 - (12,500-2,800)*$26

Direct material quantity variance = 9,660*$26 - 9,700*$26

Direct material quantity variance = $251,160 - 252,200 = $1,040 Unfavorable

Company should sign the contract with new supplier as material price variance is favorable.

2.

Direct labor rate variance = Actual hours*standard rate - Actual hours*actual rate

Direct labor rate variance = 3,150*$14 - 3,150*$12

Direct labor rate variance = $44,100 - 37,800 = $6,300 Favorable

Direct labor efficiency variance = Standard hours*standard rate - Actual hours*standard rate

Direct labor efficiency variance = (4,200*0.50)*$14 - 3,150*$14

Direct labor efficiency variance = $29,400 - 44,100 = $14,700 Unfavorable

No, the new labor mix should not be continued as though labor rate variance is favorable , labor efficiency variance is unfavorable which results in unfavorable total labor cost variance

3.

Variable overhead rate variance = Actual hours*standard rate - Actual hours*actual rate

Variable overhead rate variance = 3,150*$3.40 - $4,200

Variable overhead rate variance = $10,710 - 4,200 = $6,510 Favorable

Variable overhead efficiency variance = Standard hours*standard rate - Actual hours*standard rate

Variable overhead efficiency variance = (4,200*0.5)*$3.40 - 3,150*$3.40

Variable overhead efficiency variance = $7,140 - 10,710 = $3,570 Unfavorable